# AI Flow at the Network Edge

### v3

@@ I Introduction
We are witnessing a transformative era in the field of artificial intelligence (AI) with groundbreaking technologies [1, 2, 3]. LLMs, like ChatGPT and PaLM, have demonstrated remarkable capabilities in natural language understanding, generation, and reasoning. Going beyond language, multimodal foundation models, like CLIP and Llama, have showcased exceptional performance in cross-modal perception, understanding, and interaction tasks. These advancements led to their adaptations in a broad spectrum of application domains, including embodied robotics, augmented reality, and autonomous driving [4, 5, 6, 7]. Most recently, edge devices (mobile phones, smart wearables, and IoT sensors) are becoming increasingly widespread, and sensory data are easy to access. This has sparked a surge of interest in pushing intelligence applications from the central cloud to the network edge [8, 9]. As visualized in Fig. 1, the future envisions a scenario where AI technologies are seamlessly integrated into various aspects of daily life.
@@
While offering exciting opportunities, deploying large models at the network edge faces new challenges. In contrast to cloud intelligence, performing AI tasks at the network edge differs significantly in many cases. First, the fundamental difference is the available resources. Edge devices are usually equipped with limited computation resources, whereas cloud servers possess a large number of powerful processing units. As a result, on-device model inference struggles to support real-time responses since the large foundation models demand intensive computation [10]. Besides, edge devices are often wirelessly connected. The limited uplink bandwidth and the highly dynamic nature of wireless channels hinder transmitting large volumes of data collected by edge devices to cloud servers in real-time.
Aiming at efficient delivery of diversified intelligent services, we propose AI Flow, a framework that seamlessly integrates intelligence capabilities directly at the network edge. Specifically, AI Flow aims to streamline the inference process by jointly leveraging the heterogeneous computational resources available across devices, edge nodes, and cloud servers. This framework distributes the inference workload across different network layers and adapts to dynamic network conditions. To facilitate cooperation among multiple computational nodes, AI Flow provides a paradigm shift in the design of communication network systems from transmitting information flow to intelligence flow. The goal of communications needs to be folded into the inference process. Instead of sending raw data from edge devices to servers for processing, the intelligence flow features a task-oriented property, where edge devices extract only critical features from the raw sensory data and discard redundant information to reduce communication overhead. To summarize, the advantages of the AI Flow framework are manifold.
@@ III-C Key Design Considerations
2) How to speed up model inference? The most notable characteristic of large models is the massive number of parameters, bringing significant computational demands. The computational complexity of transformer-decoder models increases quadratically with the token length. While methods like key-value (KV) cache mitigate this challenge by avoiding recomputation, the substantial memory consumption bottlenecks model inference.
@@ IV-B Enabling Techniques
The synergy of large and small models can also support cooperative inference. As illustrated in Fig. 3, one such approach is speculative decoding [13]. Typical LLMs and VLMs need to generate tokens in an autoregressive manner that outputs tokens iteratively. Speculative decoding introduces a novel inference paradigm where a small model deployed at the edge device drafts multiple tokens once. Then a large model at the server verifies and corrects these tokens in parallel. When the acceptance rate of draft tokens by the large model is high, significant latency reduction can be achieved compared to the traditional server-only inference. This is due to the low cost of generating draft tokens by a small model, coupled with the low overhead of transmitting these tokens.
@@ V-B Enabling Techniques
There are many empirical successes in speeding up model inference through compression techniques. Model parameter pruning is devoted to removing unimportant components [15]. By aggregating useful information into a subset of activations, other neurons or channels containing task-irrelevant information can be pruned without performance loss. Similarly, low-rank factorization decomposes parameter matrices into lower-dimensional components, eliminating redundant information and reducing the model size. Neural architecture search leverages reinforcement learning to design model architectures that balance performance with computational efficiency. Besides, model quantization decreases the informativeness of activations by transforming high-precision weights into lower-bit float or integer values. Alternatively, knowledge distillation transfers knowledge from a large and complex model to a compact one to mimic the functionality of the teacher model, yet with fewer parameters and often faster inference times.
@@
Furthermore, there are complementary techniques that can enhance the inference efficiency of large models. During the autoregressive generation, KV pairs from previous tokens are stored in a cache to assist in generating future tokens. However, given a long input sequence, the cache size increases rapidly, leading to long inference latency. Popular optimization techniques include quantizing the KVs and evicting the less informative KVs. These approaches allow for faster token generation by strategically compressing the cache.
@@ VI Case Study: Speculative Decoding for Edge Inference
This section provides performance evaluations to demonstrate the performance of AI Flow, which leverages the speculative decoding approach to enhance the efficiency of cooperative inference. Consider a device-edge cooperative inference system. The edge device is equipped with a GeForce RTX 4090, and the edge node is a powerful GPU cluster. They are connected by wireless channels with data rates ranging from 500 KB/s to 2 MB/s. We evaluate the performance of speculative decoding on an image captioning task, as shown in Fig. 5. Due to the limitation of local resources, a small InternVL2-2B model is deployed on the device to generate draft tokens. Meanwhile, a large InternVL2-26B model, which serves as the target model, is deployed on the edge server. For the baseline method, known as server-only inference, the edge device transmits raw images directly to the edge server. The server then uses the target model for inference. The experiments are conducted on the Vehicles-OpenImage dataset, which is a slice of the OpenImage dataset containing 627 images of various vehicle classes. We focus on the response latency in the experiment and select the time per output token (TPOP) as the metric. For a fair comparison, we maintain that the outputs from speculative decoding have the same distributions as those from the server-only inference.
@@
As shown in Fig. 6, speculative decoding significantly reduces the TPOP compared with the baseline method. In particular, the optimal draft token length in this experiment is 4, where the inference speed is approximately double that of the baseline. This optimal length balances the acceptance rate of draft tokens with their quantity, maximizing the number of tokens that pass verification at the edge server. When the draft token length is set below 4, there is a slight increase in response latency. This occurs because generating fewer draft tokens does not fully leverage the predictive capabilities of the smaller model on the edge device. In contrast, setting the draft token length above 4 leads to an increase in response latency. This is because the smaller model on the edge device produces more errors with longer token sequences, resulting in a lower rate of token acceptance. Specifically, once the token length exceeds 22, the overall latency is higher than that of the baseline. Therefore, carefully selecting the optimal draft token length has the potential to enhance the performance gains of speculative decoding in cooperative inference systems.

### v4

@@ I Introduction
We are witnessing a transformative era in the field of artificial intelligence (AI) with groundbreaking technologies [1, 2]. LLMs, like ChatGPT and PaLM, have demonstrated remarkable capabilities in natural language understanding, generation, and reasoning. Going beyond language, multimodal foundation models, like CLIP and Llama, have showcased exceptional performance in cross-modal perception, understanding, and interaction tasks. These advancements led to their adaptations in a broad spectrum of application domains, including embodied robotics, augmented reality, and autonomous driving [3, 4, 5, 6]. Most recently, edge devices (mobile phones, smart wearables, and IoT sensors) are becoming increasingly widespread, and sensory data are easy to access. This has sparked a surge of interest in pushing intelligence applications from the central cloud to the network edge [7, 8]. As visualized in Fig. 1, the future envisions a scenario where AI technologies are seamlessly integrated into various aspects of daily life.
@@
While offering exciting opportunities, deploying large models at the network edge faces new challenges. In contrast to cloud intelligence, performing AI tasks at the network edge differs significantly in many cases. First, the fundamental difference is the available resources. Edge devices are usually equipped with limited computation resources, whereas cloud servers possess a large number of powerful processing units. As a result, on-device model inference struggles to support real-time responses since the large foundation models demand intensive computation [9]. Besides, edge devices are often wirelessly connected. The limited uplink bandwidth and the highly dynamic nature of wireless channels hinder transmitting large volumes of data collected by edge devices to cloud servers in real time.
Existing studies on edge AI typically focus on small models while overlooking the complexities and resource demands of large-scale transformer-based models. To address this gap and efficiently provide diverse intelligent services, we propose AI Flow, a framework that optimizes the deployment and execution of both small and large models. Specifically, AI Flow streamlines the inference process by jointly leveraging the heterogeneous computational resources available across devices, edge nodes, and cloud servers. This framework distributes the inference workload across different network layers and adapts to dynamic network conditions. To facilitate cooperation among multiple computational nodes, AI Flow provides a paradigm shift in the design of communication network systems from transmitting information flow to intelligence flow. The goal of communications needs to be folded into the inference process. Instead of sending raw data from edge devices to servers for processing, the intelligence flow features a task-oriented property, where edge devices extract only critical features from the raw sensory data and discard redundant information to reduce communication overhead. To summarize, the advantages of the AI Flow framework are manifold.
@@ III-C Key Design Considerations
2) How to speed up model inference? The most notable characteristic of large models is the massive number of parameters, bringing significant computational demands. The computational complexity of transformer-decoder models increases quadratically with the token length. While methods like key-value (KV) cache mitigate this challenge by avoiding recomputation, substantial memory consumption limits the scalability. This results in frequent memory overflows and creates new challenges for memory management. In addition, real-time inference in large models is often slow due to their autoregressive nature.
@@ IV-B Enabling Techniques
The synergy of large and small models can also support cooperative inference. As illustrated in Fig. 3, one such approach is speculative decoding [12]. Typical LLMs and VLMs need to generate tokens in an autoregressive manner that outputs tokens iteratively. This process is slow and computationally intensive. Speculative decoding facilitates the simultaneous decoding of multiple tokens. It introduces a novel inference paradigm where a small model deployed at the edge device drafts multiple tokens once. Then a large model at the server verifies and corrects these tokens in parallel. When the acceptance rate of draft tokens by the large model is high, significant latency reduction can be achieved compared to the traditional server-only inference. This is due to the low cost of generating draft tokens by a small model, coupled with the low overhead of transmitting these tokens.
@@ V-B Enabling Techniques
There are many empirical successes in speeding up model inference through compression techniques. Model parameter pruning is devoted to removing unimportant components. By aggregating useful information into a subset of activations, other neurons or channels containing task-irrelevant information can be pruned without performance loss. Similarly, low-rank factorization decomposes parameter matrices into lower-dimensional components, eliminating redundant information and reducing the model size. Neural architecture search leverages reinforcement learning to design model architectures that balance performance with computational efficiency. Besides, model quantization decreases the informativeness of activations by transforming high-precision weights into lower-bit float or integer values. Knowledge distillation trains a small student model to mimic the functionality of a large teacher model. Alternatively, the adoption of explainable AI [14] forecasts the workload and identifies critical components in the neural architecture. This enables model designers to adjust the tradeoff between model complexity and performance drop.
@@
Furthermore, there are complementary techniques that can enhance the inference efficiency of large models. During the autoregressive generation, KV pairs from previous tokens are stored in a cache to assist in generating future tokens. However, given a long input sequence, the cache size increases rapidly, leading to long inference latency. Popular optimization techniques include quantizing the KVs and evicting the less informative KVs. These approaches allow for faster token generation by strategically compressing the cache. Additionally, some novel low-complexity attention mechanisms, such as the Informer, can be adopted to accelerate inference [15].
@@ VI Case Study: Speculative Decoding for Edge Inference
This section provides performance evaluations to demonstrate the performance of AI Flow, which leverages the speculative decoding approach to enhance the efficiency of cooperative inference.
VI-A Setup
Consider a device-edge cooperative inference system. The edge device is equipped with a GeForce RTX 4090, and the edge node is a powerful GPU cluster. They are connected by wireless channels with data rates ranging from 500 KB/s to 2 MB/s. We evaluate the performance of speculative decoding on an image captioning task served at the edge, as shown in Fig. 5. The experiments are conducted on the Vehicles-OpenImage dataset, which is a subset of the OpenImage dataset containing 627 images of various vehicle classes in open space. The small and large models are selected from the InternVL2 series, which is a high-scalable VLM family that includes models ranging from a 1B model to a significantly more powerful 108B model. In the experiment, a small InternVL2-2B model is deployed at the device to generate draft tokens. Meanwhile, a large InternVL2-26B model is deployed on the edge server. For the baseline method, known as server-only inference, the edge device transmits raw images directly to the edge server. The server then uses the large model for inference. We focus on the response latency in the experiment and select the time per output token (TPOT) as the metric. To maintain a fair comparison, the large model in speculative decoding corrects all inconsistent tokens generated by the small model. This guarantees that the final output of speculative decoding matches exactly with the output of server-only inference.
@@ VI-A Setup
VI-B Results
As shown in Fig. 6, speculative decoding significantly reduces the TPOT compared with the baseline method. In particular, the optimal draft token length in this experiment is 4, where the inference speed is approximately double that of the baseline. This optimal length balances the acceptance rate of draft tokens with their quantity, maximizing the number of tokens that pass verification at the edge server. When the draft token length is set below 4, there is a slight increase in response latency. This occurs because generating fewer draft tokens does not fully leverage the predictive capabilities of the smaller model on the edge device. In contrast, setting the draft token length above 4 leads to an increase in response latency. This is because the smaller model on the edge device produces more errors with longer token sequences, resulting in a lower rate of token acceptance. Specifically, once the token length exceeds 22, the overall latency is higher than that of the baseline. Therefore, carefully selecting the optimal draft token length has the potential to enhance the performance gains of speculative decoding in cooperative inference systems.
VI-C Discussion
In this case study, we focus on the effectiveness and efficiency of the speculative decoding method for cooperative inference. Empirical results show that the device-edge cooperative scheme can largely reduce response latency compared to server-only inference. It is important to note that, as discussed in Section V, many other inference speedup techniques, such as model compression, dynamic neural networks, and KV cache optimization, are available. These techniques are complementary to the cooperative inference scheme and can be combined with speculative decoding to achieve even greater efficiency. For instance, model compression could further reduce the computational load on both the device and the server, while KV cache optimization could significantly improve the speed of token verification. Exploring the interplay between these techniques represents a promising direction for future research. By combining multiple acceleration strategies, it is possible to achieve significant performance gains in edge inference scenarios.